OpenStudy (anonymous):

The question is Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). please help

OpenStudy (anonymous):

This is algebra 1

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

First, solve for Y. Y = -3X+7 Now you have the slope, -3. Perpendicular, you need the opposite reciprocal of the slope. \[Y =\frac{ 1 }{ 3 }X + 7\]

OpenStudy (anonymous):

Now you have the slope, take off the +7 and Plug in the points

OpenStudy (anonymous):

-1 = (1/3)*6 + b -1 = 2 + b. solve for b.

OpenStudy (anonymous):

Do you get it now?

OpenStudy (anonymous):

yeah and b is the y intercept right

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

thanks for the help